# A Hypernetwork-Based Approach to KAN Representation of Audio Signals

### v2

@@ Abstract.
conference: ; isbn: ;ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Learning paradigmsccs: Applied computing Sound and music computing
Figure 1. Sound representation is a challenging task. In classical analysis, we usually apply the FFT (Fast Fourier Transform), which analyzes frequencies by separating them into sinusoidal components. We propose to analyze sounds using KAN (Kolmogorov-Arnold Network), which uses trainable activation functions instead of classical layers. Such an operation can be seen as approximating signals by a linear combination of basis functions, bearing a similarity to the FFT or wavelets.
@@ 1. Introduction
Implicit neural representations (INRs) depict discrete data in neural network weights. INR finds many applications in the case of image (Klocek et al., 2019), video (Chen et al., 2022), or 3D objects (Spurek et al., 2020). In the case of sound, INRs are unexplored, and only a few approaches are dedicated to such data. HyperSound (Szatkowski et al., 2023) presents an autoencoder-based architecture that takes raw input sound and produces weight for INR-based sound representation. INRAS (Su et al., 2022) is an INR network that, by mapping the scene coordinates to the corresponding impulse responses, enables the modeling of spatial audio of interior scenes. Siamese SIREN (Lanzendörfer and Wattenhofer, 2023), is the extension of SIREN (Sitzmann et al., 2020) used for audio compression.
@@
The recently proposed KAN (Kolmogorov-Arnold Network) (Liu et al., 2024a) replaces traditional fully connected layers with activation functions. While KAN shows potential, it faces limitations with high-dimensional data. However, its effectiveness in low-dimensional settings and successful application to complex function regression tasks make it a promising candidate for audio INR.
@@
This paper shows that KAN is effective in modeling individual audio signals and can be extended to various functions through the hypernetwork paradigm. For single audio INR, KAN trains more rapidly and achieves better results. Setting up the conditions for a KAN is not immediately straightforward. We show that hypernetworks can generate KANs. In practice, we can directly produce KAN parameters or use updates for the universal KAN-based architecture. Hence, we propose two methods, a classical HyperSound (Szatkowski et al., 2023) with a KAN target network. The second uses a few-shot setting (Tancik et al., 2021; Batorski et al., 2024) in INR, which is an implementation of the Model-Agnostic Meta-Learning algorithm (MAML) (Finn et al., 2017). As there is no few-shot model for audio signals, we introduce FewSound  [^1] The source code for the proposed method can be found at the following link: https://github.com/gmum/fewsound. It takes raw sound on input and uses a hypernetwork to generate updates to the universal weights of INR dedicated to encoding sound. In our work, we examine six approaches: SIREN (Sitzmann et al., 2020), NeRF (Mildenhall et al., 2020), Random Fourier Features (Tancik et al., 2020), WIRE (Saragadam et al., 2023), FINER (Liu et al., 2024b) and KAN (Liu et al., 2024a). In the model, we simultaneously train the hypernetwork weights and the universal weights. Thus, two settings show that KAN can be effectively modeled by hypernetworks.
Our study demonstrates that KAN can be viewed as INR of sound. Such models yield favorable outcomes when dealing with individual audio signals and excel in large Hypernetwork-based models trained on large datasets. We use KANs due to their spline-based activations. Since any function f (e.g., an audio representation of the magnitude over time) can be approximated by splines with $\sup\|f\_spline\|=\mathrm{O}(h^{4})$ (Hall and Meyer, 1976), KANs inherit this theoretical precision as a linear combination of splines, resulting in a polynomial bound. This suggests that KANs could be effective for audio reconstruction.
@@
  • introducing FewSound, a hypernetwork-based meta-learning method, which is used with KAN network and improves over previous hypernetwork approach HyperSound (Szatkowski et al., 2023),
@@ 2. Related Work
Initially, the INRs used ReLU-based MLPs. However, due to the ReLU’s second derivative being zero, they cannot capture the fine details of the information located in higher-order derivatives leading to worse performance. To counteract this, methods such as SIREN (Sitzmann et al., 2020) have been developed, which use a periodic activation function the sine in place of ReLU. The i-th layer of the network is defined as: $\phi_{i}(x_{i})=\sin(\omega_{i}W_{i}^{T}x_{i}+b_{i}),$ where $x_{i}$ is layer’s input, while $W_{i}$ and $b_{i}$ are, respectively, weights and biases. The use of the sine function enables SIREN to capture the details of the processed signals to a much greater degree than the widely used activation functions. Another example of a network architecture using a specialized activation function is WIRE (wavelet implicit neural representation) (Saragadam et al., 2023). Inspired by harmonic analysis, WIRE uses complex Gabor wavelet $\psi$ as its nonlinearity activation: $\psi(x)=e^{i\omega_{0}x}e^{-\lvert s_{0}x\rvert^{2}}$.
INRs can suffer from a phenomenon called spectral bias (Rahaman et al., 2019) – the preference of MLPs to learn low-frequency functions and ignore high-frequency noise, which can lead to performance degradation when high-frequency components of the input carry meaningful information (Gorji et al., 2023). A common way to overcome spectral bias is by introducing high-frequency embeddings (Mildenhall et al., 2020; Tancik et al., 2020) that modify the space over which the MLP operates, altering the frequency of its output function. FINER (Liu et al., 2024b) is an MLP-based architecture designed to flexibly tune the spectral bias. It uses a variable-periodic activation $a(x)=sin(\omega_{0}\lvert x+1\rvert x)$ and a specialized bias initialization schema in its first layer.
NeRF (Neural Radiance Field) (Mildenhall et al., 2020) models scenes via a 5D input—spatial location $(x,\,y,\,z)$ and viewing direction $(\theta,\,\phi)$—with an MLP outputting color $c=(r,\,g,\,b)$ and density $\sigma$, enabling 3D reconstruction from sparse views.
Figure 2. The overview of FewSound architecture. The input row sound is processed by encoding network $E(\cdot)$ and delivered to the fully connected Hypernetwork $H(\cdot)$ together universal weights. The Hypernetwork transforms them and returns the update of weights $\Delta\theta$ for the target INR $f_{\theta^{\prime}}$. The target network $f_{\theta^{\prime}}$ transforms time samples into the amplitude of the sound, dedicated to the considered task.
@@
The audio encoders we rely on are classified into two main types: neural speech codecs and those employed in automatic speech recognition (ASR) or ASR-adjacent solutions.
@@
KANs (Liu et al., 2024a) is a type of neural network architecture used as an alternative to MLPs, instead of using fixed activation functions. KANs utilize learnable activation functions placed on edges rather than nodes. KAN’s learnable functions are parametrized as a spline and serve as ”weights”, whereas the nodes only sum the incoming signals. This is described by the following formula:
@@
where $\Phi_{q}$ is a KAN layer, $\phi_{q,\,p}$ is a single activation function and $x_{p}$ is input feature. The authors also introduced an architecture of LANs (learnable activation networks) – the only difference between them and MLP lies in the learning aspect of the activation functions. The authors show that a network can be used in implicit neural representation tasks.
@@ 3. KAN in hypernetwork-based architectures
In this section, we show how KAN works with hypernetworks. We use two main approaches. First, we use the autoencoder-based solution HyperSound (Szatkowski et al., 2023), where we add KAN as the target model. We also examine KAN in a few-shot setting introduced by (Wang et al., 2020). To examine such a model, we propose FewSound, which is a realization of (Wang et al., 2020) in the case of KAN and hypernetworks.
@@
In the foundational study (Ha et al., 2016) hyper-networks are described as neural models that produce weights for another network tasked with solving a specific problem. The researchers aim to decrease the number of trainable parameters by creating a hyper-network with fewer parameters than the target network. Drawing a parallel between hypernetworks and generative models, the authors of (Sheikh et al., 2017) employ this technique to generate a variety of target networks that approximate the same function.
@@
A significant limitation of deep learning models is their reliance on extensive datasets. In contrast, humans can learn effectively from just a few examples. A few-shot learning (Wang et al., 2020) approach was proposed to bridge the gap between deep neural network training and human learning capabilities. This strategy treats the data set as consisting of numerous small tasks. Few-shot architectures can be quickly adapted to solve new problems in a single step. The most popular approach to few-shot learning is MAML (Finn et al., 2017). MAML uses a set of universal weights, fine-tuned for specific small tasks through a limited number of gradient updates. Due to the model-agnostic nature of the approach, it can be adapted to various models and corresponding learning algorithms.
The terminology of few-shot learning is often inconsistent due to varying definitions in existing research. For a standardized framework, see  (Chen et al., 2019; Wang et al., 2020). However, such a nomenclature is dedicated to a classification, regression, or generative model task. In the case of the INR-based model, we refer to (Tancik et al., 2021). We have a data set $X$ that contains sound samples, and our goal is to produce INR $f_{\theta}$ that encodes sound in network weights.
@@
Now we are ready to present FewSound – a model that uses Hypernetworks to update the INR weights dedicated to sound. The core concept of this update mechanism is to use information extracted from raw sound data and universal weights to determine the optimal updates for a particular task. This method adjusts the INR’s parameters for entirely different sounds.
The schema of FewSound is provided in Fig. 2. We aim to reconstruct the training example using the target network. In practice, we run our INR on all steps to reconstruct the input.
Following our previous analysis of INR in the sound reconstruction task, we consider the parameterized function $f_{\theta}$ to be one of the discussed INRs. In addition, we distinguish the trainable encoding network $E(\cdot)$ in our architecture, which transforms raw audio data into low-dimensional representation. We investigated various encoder architectures and report their performance in Tab. 5.
The encoded sound is concatenated with universal weights propagated by the universal weight encoder $\mathbf{G}(\cdot)$ and then to the fully connected hypernetwork $H(\cdot)$. The hypernetwork then modifies the universal weights: $\theta^{\prime}=\theta+\Delta_{\theta}.$ In contrast to the traditional MAML approach, which uses a gradient-based adaptation step, we propose employing a Hypernetwork to anticipate the values directly from the data.
Target model parameters are computed via Eq. 1.
@@
While FewSound accepts a fixed-size input (in our case 32768), it can be easily used on audio samples of arbitrary length by reconstructing them piece by piece and smoothing during merging.
@@
For training the model, we assume that the encoder $E(\cdot)$ is parametrized by $\gamma$, $E:=E_{\gamma}$. The universal weight encoder $G(\cdot)$ is parameterized by $\delta$, and the Hypernetwork $H(\cdot)$ by $\eta$, $H:=H_{\eta}$. During training, we first sample the batch of sounds $\mathcal{B}$ from the dataset $X$. Next, for each task, we calculate the update $\Delta_{\theta_{i}}$ using the hypernetwork and provide the updated parameters $\theta^{\prime}$ according to the rule given in (1). Finally, the objective to train the parameters of the system is calculated using the timesteps from [1, 1]11[-1,\,1][ - 1 , 1 ]:
(2) $\mathcal{L}_{HyperMAML}(f_{\theta})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta+\Delta_{\theta}})~{},$
where $\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\lambda_{t}L_{t}(x,\hat{x})+\lambda_{f}L_{f}(x,\hat{x})$, with $L_{t}$ is L1 loss in temporal space, $L_{f}$ is multi-resolution mel-scale STFT loss in frequency space (Yamamoto et al., 2020) and $\lambda$s are coefficients controlling influence of particular losses. The Enconder parameters, universal weight encoder, Hypernetwork, and global parameters $\theta$ represent the meta parameters of the system and are updated with stochastic gradient descent (SGD) by optimizing $\mathcal{L}_{HyperMAML}(f_{\theta})$.
@@
[…]
@@
[…]
@@ 4.1. Implicit Neural Representations
We compare the aforementioned INR architectures using two subsets of the VCTK (Veaux et al., 2013) dataset. Each of the subsets is generated by randomly choosing one recording per speaker, resampling it to 22.05 kHz and randomly cropping it to a desired length. We crop the recordings to 32,768 samples (about 1.5 seconds) in the first subset and to 65,536 (about 3 seconds) in the second subset. Both subsets consist of a total of 110 recordings. To make a fair comparison, all of the MLP based INRs have the same number of layers and with the same sizes, while KANs have their architecture chosen in a way that the number of parameters is similar to the other models. All architectures are trained for 10,000 steps, using the AdamW optimizer (Loshchilov and Hutter, 2019) and the same loss function as in FewSound. Table 1 shows the results of the experiments. One may note that, with the exception of MSE, the shorter audios result in better quality. KAN is competitive for most metrics, achieving the best mean score in LSD, PESQ, STOI and CDPAM and second best PSNR for 1.5 s audio; as well as the best mean LSD, PESQ, STOI, and second best MSE, PSNR, SI-SNR, CDPAM for 3 s audio.

### v3

# As Good as It KAN Get: High-Fidelity Audio Representation

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761405isbn: 979-8-4007-2040-6/2025/11ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Learning paradigmsccs: Applied computing Sound and music computing
Figure 1. Sound representation is a challenging task. In classical analysis, we usually apply the FFT (Fast Fourier Transform), which analyzes frequencies by separating them into sinusoidal components. We propose to analyze sounds using KAN (Kolmogorov-Arnold Network), which uses trainable activation functions instead of classical layers. Such an operation can be seen as approximating signals by a linear combination of basis functions, bearing a similarity to the FFT or wavelets.:
@@ 1. Introduction
Implicit neural representations (INRs) depict discrete data in neural network weights. INR finds many applications in the case of image (Klocek et al., 2019), video (Chen et al., 2022), or 3D objects (Spurek et al., 2020). In the case of sound, INRs are unexplored, and only a few approaches are dedicated to such data. HyperSound (Szatkowski et al., 2023) presents an autoencoder-based architecture that takes raw input sound, and produces weight for INR-based sound representation. INRAS (Su et al., 2022) is an INR network that, by mapping the scene coordinates to the corresponding impulse responses, enables the modeling of spatial audio of interior scenes. Siamese SIREN (Lanzendörfer and Wattenhofer, 2023) is the extension of SIREN (Sitzmann et al., 2020) used for audio compression.
@@
The recently proposed KAN (Kolmogorov-Arnold Network) (Liu et al., 2024a) replaces traditional fully connected layers with learnable activation functions based on splines. Any function f (e.g., an audio representation of the magnitude over time) can be approximated by splines with $\sup\|\text{f-spline}\|=\mathrm{O}(h^{4})$ (Hall and Meyer, 1976). KANs inherit this theoretical precision as a linear combination of splines, resulting in a polynomial bound. This indicates potential effectiveness for audio reconstruction. While KAN shows potential, it faces limitations with high-dimensional data. However, its effectiveness in low-dimensional settings and successful application to complex function regression tasks make it a promising candidate for audio INR.
@@
In our work we examine six INR approaches: SIREN (Sitzmann et al., 2020), NeRF (Mildenhall et al., 2020), Random Fourier Features (Tancik et al., 2020), WIRE (Saragadam et al., 2023), FINER (Liu et al., 2024b) and KAN (Liu et al., 2024a). This paper shows that KAN is effective in modeling individual audio signals and can be extended to various functions through the hypernetwork paradigm. For single audio INR, KAN trains more rapidly and achieves better results. We also show that KANs can be generated by hypernetworks. In practice, we can directly produce KAN parameters or use updates for a universal KAN-based architecture. Hence, we propose two methods. The first one is classical HyperSound (Szatkowski et al., 2023) with a KAN target network, while the second one uses a few-shot setting (Tancik et al., 2021; Batorski et al., 2024) in INR. As there is no few-shot model for audio signals, we introduce FewSound.  [^1] The source code for the proposed method can be found at the following link: https://github.com/gmum/fewsound It takes raw sound on input and uses a hypernetwork to generate updates to the universal weights of an INR dedicated to encoding sound. FewSound is an implementation of the Model-Agnostic Meta-Learning algorithm (MAML) (Finn et al., 2017). Thus, two settings show that KAN can be effectively modeled by hypernetworks.
@@
  • introducing FewSound, a hypernetwork-based meta-learning method, which is used with KAN networks and improves over previous hypernetwork approach HyperSound (Szatkowski et al., 2023),
@@ 2. Related Work
Initially, the INRs used ReLU-based MLPs. However, due to the ReLU’s second derivative being zero, they cannot capture the fine details of the information located in higher-order derivatives leading to worse performance. To counteract this, methods as SIREN (Sitzmann et al., 2020) have been developed, which use a periodic activation function in place of ReLU. The i-th layer of the network is defined as: $\phi_{i}(x_{i})=\sin(\omega_{i}W_{i}^{T}x_{i}+b_{i}),$ where $x_{i}$ is layer’s input, while $W_{i}$ and $b_{i}$ are, respectively, weights and biases. The use of the sine function enables SIREN to capture the details of the processed signals to a much greater degree than the widely used activation functions. Another example of a network architecture using a specialized activation function is WIRE (wavelet implicit neural representation) (Saragadam et al., 2023). Inspired by harmonic analysis, WIRE uses complex Gabor wavelet $\psi$ as its nonlinearity activation: $\psi(x)=e^{i\omega_{0}x}e^{-\lvert s_{0}x\rvert^{2}}$.
INRs are prone to spectral bias (Rahaman et al., 2019), where MLPs favor low-frequency functions and neglect high-frequency signals, reducing performance when such components are informative (Gorji et al., 2023). This is commonly mitigated with high-frequency embeddings (Mildenhall et al., 2020; Tancik et al., 2020), which reshape the input space and adjust output frequencies.
FINER (Liu et al., 2024b), an MLP-based architecture, flexibly tunes spectral bias via a variable-periodic activation $a(x)=\sin(\omega_{0}\lvert x+1\rvert x)$ and specialized first-layer bias initialization.
NeRF (Neural Radiance Field) (Mildenhall et al., 2020) represents scenes with a 5D input: spatial location (x,y,z)(x,\,y,\,z) and viewing direction $(\theta,\,\phi)$, and an MLP that outputs color c=(r,g,b)c=(r,\,g,\,b) and density $\sigma$, enabling 3D reconstruction from sparse views.
@@
In our research, we use learned representations from deep neural networks, which outperform traditional hand-crafted features (Baevski et al., 2020). These representations fall into 2 categories: acoustic embeddings from neural speech codecs (Défossez et al., 2023), and semantic embeddings (Hsu et al., 2021). We focus on representations derived from models designed for ASR or related tasks.
@@
Figure 2. The overview of FewSound architecture. The input raw sound is processed by encoding network $E(\cdot)$ and, together with encoded universal weights, passed to the fully connected Hypernetwork $H(\cdot)$. The Hypernetwork transforms them and returns the update of weights $\Delta\theta$ for the target INR $f_{\theta^{\prime}}$. Finally, the target network $f_{\theta^{\prime}}$ maps time samples to sound amplitudes specific to the considered task. :
@@
KAN (Liu et al., 2024a) is a type of neural network architecture used as an alternative to MLPs, instead of using fixed activation functions. KANs utilize learnable activation functions placed on edges rather than nodes. KANs’ learnable functions are parametrized as a spline and serve as ”weights”, whereas the nodes only sum the incoming signals. This is described by the following formula:
@@
where $\Phi_{q}$ is a KAN layer, $\phi_{q,\,p}$ is a single activation function and $x_{p}$ is input feature. The authors introduced LANs (learnable activation networks), which differ from MLPs only by learning activation functions. They demonstrate that such networks can be applied to implicit neural representation tasks.
@@ 3. KAN in hypernetwork-based architectures
In this section, we demonstrate how KAN integrates with hypernetworks using two main approaches. First, we extend the autoencoder-based HyperSound (Szatkowski et al., 2023) by employing KAN as the target model. Second, we explore KAN in a few-shot setting introduced by (Wang et al., 2020). To this end, we propose FewSound, which adapts the framework of (Wang et al., 2020) to the context of KANs and hypernetworks.
@@
In the seminal work (Ha et al., 2016), Hypernetworks are introduced as neural models that produce the weights of another network dedicated to solving a specific task. The approach aims to reduce the number of trainable parameters by employing a hypernetwork with fewer parameters than the target network. Building on this concept, numerous studies have emerged. Among them, (Sheikh et al., 2017) draws an analogy to generative models, demonstrating how this technique can produce a variety of target networks approximating the same function. In the audio INR domain, HyperSound (Szatkowski et al., 2023) leverages hypernetworks to transform raw audio into weights of an INR-based representation.
@@
A key limitation of deep learning models is their dependence on large datasets, whereas humans can learn effectively from just a few examples. Few-shot learning (Wang et al., 2020) seeks to bridge this gap by structuring data into numerous small tasks and using architectures capable of rapid adaptation to new problems in a single step. The most popular method, MAML (Finn et al., 2017), leverages a set of universal weights that are fine-tuned for specific small tasks through a limited number of gradient updates. Its model-agnostic nature allows it to be applied to a wide range of models and learning strategies. Nevertheless, further generalization is possible. HyperMAML (Przewięźlikowski et al., 2024) achieves this by replacing the gradient-based adaptation step with a Hypernetwork that predicts task-specific weight updates directly from the data.
The terminology of few-shot learning is often inconsistent due to differing definitions in prior work. Standardized frameworks exist (Chen et al., 2019; Wang et al., 2020), but their nomenclature is focused on a classification, regression, or generative task. For INR-based models, we follow (Tancik et al., 2021). In this setting, we have a dataset $X$ of sound samples and aim to produce an INR $f_{\theta}$ that encodes the audio signals in its weights.
@@
At this stage, we are ready to present FewSound, a model that employs Hypernetworks to adapt universal INR weights specialized for sound. As illustrated in Fig. 2, its key idea lies in integrating the information extracted from raw sound data and universal weights to determine the optimal adjustment for a particular task. As a result, our method is able to generate INR’s parameters for entirely different sounds.
Within the FewSound framework, we distinguish a trainable encoding network $E(\cdot)$ that maps raw audio data to a low-dimensional representation $\mathbf{E}_{\mathcal{S}}$. We explore several encoder architectures, including SoundStream, Whisper, SNAC, and Spectrogram Encoder, and report their performance in Tab. 5. In addition, we define a universal weight encoder $\mathbf{G}(\cdot)$ that compresses universal weights $\theta$ to a representation $\mathbf{E}_{\mathcal{\theta}}$ of the same dimensionality as $\mathbf{E}_{\mathcal{S}}$. Both representations are concatenated and subsequently processed by a fully connected hypernetwork $H(\cdot)$, which outputs the adjustment $\Delta_{\theta}$ to the universal weights for the given input task. Consequently, the final target model parameters $\theta^{\prime}$ can be expressed as:
@@
Using these adapted weights, we can reconstruct the training example by running one of the INRs described in Section 2 over all time indices of the input signal. Although FewSound accepts a fixed-size input (in our case 32768), it can also be used on audio samples of arbitrary length by reconstructing them piece by piece in an overlapping window scheme and applying smoothing during concatenation.
@@
In the training procedure, we assume that the encoder $E(\cdot)$ is parameterized by $\gamma$ ($E:=E_{\gamma}$), the universal weight encoder $G(\cdot)$ by $\delta$, and the Hypernetwork $H(\cdot)$ by $\eta$ ($H:=H_{\eta}$). During each training iteration, we first sample the batch of sounds $\mathcal{B}$ from the dataset $X$. Next, the hypernetwork computes the task-specific update $\Delta_{\theta_{i}}$, yielding updated weights $\theta^{\prime}_{i}$ as given in (1). These parameters define the corresponding INR $f_{\theta^{\prime}_{i}}$. Finally, we obtain the reconstruction $\hat{x}$ of the input sound $x$ using time steps from [1,,1][-1,,1], and train the system parameters with the loss function:
(2) $\mathcal{L}_{FewSound{}}(f_{\theta})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}(f_{\theta^{\prime}_{i}})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}(f_{\theta+\Delta_{\theta_{i}}})~,$
[…]
@@
$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@
$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@ 4.1. Implicit Neural Representations
We compare the aforementioned INR architectures using two subsets of the VCTK (Veaux et al., 2013) dataset. Each of the subsets is generated by randomly choosing one recording per speaker, resampling it to 22.05 kHz and randomly cropping it to a desired length. We crop the recordings to 32,768 samples (about 1.5 seconds) in the first subset and to 65,536 (about 3 seconds) in the second subset. Both subsets consist of a total of 110 recordings. To make a fair comparison, all of the MLP based INRs have the same number of layers with the same sizes, while KANs have their architecture chosen in a way that the number of parameters is similar to the other models. All architectures are trained for 10,000 steps, using the AdamW optimizer (Loshchilov and Hutter, 2019) and the same loss function $\mathcal{L}$ as in Eq. 2. Table 1 shows the results of the experiments. One may note that, with the exception of MSE, the shorter audios result in better quality. KAN is competitive for most metrics, achieving the best mean score in LSD, PESQ, STOI and CDPAM and second best PSNR for 1.5 s audio; as well as the best mean LSD, PESQ, STOI, and second best MSE, PSNR, SI-SNR, CDPAM for 3 s audio.